Angle G is supplementary to angle H. Find the measure of each angle if angel G = (7x + 10) degrees, and angle H = (3x) degrees. Thanks for the help!
Answer by AdolphousC(70)

If the angles are supplementary, that means the sum of the two angles should be 180.


7x + 10 + 3x = 180 Combine Link Terms
10x + 10 = 180 Subtract 10 on both sides
10x = 170 Divide by 10 on both sides
x = 17.


if x = 17 then
G = 7(17) + 10 = 119 + 10 = 129
H = 3(17) = 51


129 + 51 = 180